• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M Friend

    Log entry time 16:04:32 on June19,2010

    Entry number 327574

    keyword=Day shift summary

    Starting shift with the spectrometer DAQs for both arms down. We did a remote reboot of everything (as well as some kcoda's and startcoda's), and the left arm is running again, although the right arm ROC1 is still down.

    8:05 - Start ramping up to 70uA.

    10:49 - Called for escorted access for the moller fill.

    12:10 - Back to beam. Quick raster check on Be0, and then up to 70uA on Pb/D #1.

    13:00 - Lost readback on the temperature monitors. The operators at MCC cannot figure out the problem, so they will page someone to hopefully figure it out.

    13:37 - Go to low current to check physics spot++. Right arm runs with VDCs on. Left arm DAQ crashed, but eventually (after going back to high current) got it going again.

    14:24 - Back up to 70uA, since we recovered readback on the beamline temeperatures (by using the green reboot button in the middle room).

    Production runs (not sure if number of coulombs is correct or needs to be multiplied by 2):

    4697 - 87.8 multiplets/98.6mC

    4698 - 72.4 mult/82mC

    4699 - 82 mult/92.4mC

    4700 - 66.2 mult/74mC

    4701 - 44.2 mult/50mC

    4702 - 82.1 mult/92.8mC

    4703 - in progress at the end of the shift

    Q2 runs (with T5 trigger, VDCs on):

    Runs 27935/6968, 6969